Sir If x4/3- 1 = 0, 4/3 is power of x Then what would be the value of x is x +1 and -1 OR x =1 only pls ans it with explanation I am confused


Sir
If x4/3- 1 = 0, 4/3 is power of x
Then what would be the value of x is x +1 and -1 OR x =1 only pls ans it with explanation I am confused

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Sunil

X4/3 = X*X1/3=1

Now take X=1 then X1/3 =1 hence equation is satisfied

Now take X=-1 then X1/3 = -1, Again equation is satisfied

Hence X= -1 and 1

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..